OpenStudy (***[isuru]***):

the roots of x^2 + ax + b = 0 is K and L . the roots of x^2 + px + q = 0 is K and N. the roots of x^2 + bx+ a = 0 is T and S . the roots of x^2 + qx + p = 0 is T and R ... if a , b, p , q are real constants then prove a+ b = p + q

OpenStudy (***[isuru]***):

@ganeshie8 and @phi pls help

OpenStudy (dape):

Factor all the equations in terms of the roots: \[ x^2+ax+b=(x-K)(x-L)=0 \\ x^2+px+q=(x-K)(x-N)=0 \\ x^2+bx+a=(x-T)(x-S)=0 \\ x^2+qx+p=(x-T)(x-R)=0 \] Now subtract the second from the first, and the last from the second to last, and factor the expressions: \[ (x-K)((x-L)-(x-N))=(x-K)(N-L)=0 \\ (x-T)((x-S)-(x-R))=(x-T)(R-S)=0 \] In both expressions, one factor must be zero, and it can't be the one containing \(x\) since \(x\) is variable. Thus \[N-L=0\Leftrightarrow N=L \\ R-S=0\Leftrightarrow R=S\] Putting this into our original factorization above, we see that the first two and the last two equations are respectively equal. Specifically, the coefficients must be equal, so \(a=p\) and \(b=q\). From this it immediately follows that \[a+b=p+q\]

OpenStudy (phi):

I think we can do this without regard to the roots: x^2 + ax + b = 0 x^2 + bx+ a = 0 add these equations to get 2x^2 +(a+b)x +(a+b) = 0 true for all x similarly x^2 + px + q = 0 x^2 + qx + p = 0 2x^2 +(p+q)x + (p+q) = 0 true for all x equating zero to zero: 2x^2 +(a+b)x +(a+b) =2x^2 +(p+q)x + (p+q) (a+b)x +(a+b) =(p+q)x + (p+q) the left side (a line) coincides the line on the right side (for all x) only if the slopes and intercepts are equal. i.e. only when a+b= p+q